OpenStudy (anonymous):

how do you simplify (2cos(t))^2+(3sin(t))^2

OpenStudy (anonymous):

Cos sqaured t is 1 minus sin squared t. So sub that in and simplify i suppose?

OpenStudy (anonymous):

\[\begin{align*}(2\cos t)^2+(3\sin t)^2&=4\cos^2t+9\sin^2t\\ &=(4\cos^2t+4\sin^2t)+5\sin^2t\\ &=4(\cos^2t+\sin^2t)+5\sin^2t \end{align*}\]